1)     Recent business graduates currently employed in full-time positions were surveyed. Family backgrounds were self-classified as relatively high or low socioeconomic status. For a random sample of 16 high- socioeconomic status recent business graduates, mean total compensation was $34500 and the sample standard deviation was $8520. For an independent random sample of 9 low-socioeconomic status recent business graduates, mean total compensation was $31,499 and the sample standard deviation was $7,521.

A)   Find a 90% confidence interval for the difference between the two populations means. 

B)    Test at 1% level of significance the means of the two populations are not equal.



1
Expert's answer
2021-02-24T06:52:59-0500

A) Assume that the population variances are unequal, so then the number of degrees of freedom is computed as follows:


"df=\dfrac{(s_1^2/n_1+s_2^2/n_2)^2}{\dfrac{(s_1^2/n_1)^2}{n_1-1}+\dfrac{(s_2^2/n_2)^2}{n_2-1}}=18.56027"

The critical value for "\alpha=0.1" and "df=18.56027" degrees of freedom is "t_{1-\alpha/2, n-1}=1.7312."

Since we assume that the population variances are unequal, the standard error is computed as follows:


"se=\sqrt{\dfrac{s_1^2}{n_1}+\dfrac{s_2^2}{n_2}}"

"=\sqrt{\dfrac{(8520)^2}{16}+\dfrac{(7521)^2}{9}}=3289.673"

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x_1}-\bar{x_2}-t_c\times se, \bar{x_1}-\bar{x_2}+t_c\times se)"

"=(34500-31499-1.7312\times 3289.673,"

"34500-31499+1.7312\times 3289.673)"

"=(-2693.424, 8695.424)"

Therefore, based on the data provided, the 90% confidence interval for the difference between the population means "\mu_1-\mu_2" is "-2693.424<\mu_1-\mu_2<8695.424," which indicates that we are 90%

confident that the true difference between population means is contained by the interval "(-2693.424, 8695.424)"


B) The provided sample means are "\bar{x}_1=34500" and "\bar{x}_2=31499."

The provided sample standard deviations are "s_1=8520, s_2=7521."

The sample sizes are "n_1=16" and "n_2=9."

The following null and alternative hypotheses need to be tested:

"H_0: \mu_1=\mu_2"

"H_1: \mu_1\not=\mu_2"

This corresponds to two-tailed test, for which a t-test for two population means, with two independent samples, with unknown population standard deviations will be used.

The number of degrees of freedom are "df=18.56027," and the significance level is "\alpha=0.01."

Based on the information provided, the critical value for a two-tailed test is "t_c=2.868."

The rejection region for this two-tailed test is "R=\{t:|t|>2.868\}"

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}_1-\bar{x}_2}{\sqrt{\dfrac{s_1^2}{n_1}+\dfrac{s_2^2}{n_2}}}=\dfrac{34500-31499}{\sqrt{\dfrac{8520^2}{16}+\dfrac{7521^2}{9}}}=0.912"

Since it is observed that "t=0.912<2.868=t_c," it is then concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that the population mean "\mu_1" is different than "\mu_2," at the 0.01 significance level.

Using the P-value approach: The p-value "df=18.56027, t=0.912," "two-tailed," is "p=0.3733," and since "p=0.3733\geq0.01," it is concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that the population mean "\mu_1" is different than "\mu_2" at the 0.01 significance level.